# Can Implicit Bias Imply Adversarial Robustness?

### v1

@@ Abstract
The implicit bias of gradient-based training algorithms has been considered mostly beneficial as it leads to trained networks that often generalize well. However, Frei et al. (2023) show that such implicit bias can harm adversarial robustness. Specifically, when the data consists of clusters with small inter-cluster correlation, a shallow (two-layer) ReLU network trained by gradient flow generalizes well, but it is not robust to adversarial attacks of small radius, despite the existence of a much more robust classifier that can be explicitly constructed from a shallow network. In this paper, we extend recent analyses of neuron alignment to show that a shallow network with a polynomial ReLU activation (pReLU) trained by gradient flow not only generalizes well but is also robust to adversarial attacks. Our results highlight the importance of the interplay between data structure and architecture design in the implicit bias and robustness of trained networks.
@@ 1 Introduction
Behind the success of deep neural networks in many application domains lies their vulnerability to adversarial attacks, i.e., small and human-imperceptible perturbations to the input data. Such a phenomenon was observed in the seminal paper of Szegedy et al. (2014) and has motivated a large body of work on building defenses against such attacks (Shafahi et al., 2019; Papernot et al., 2016; Wong et al., 2019; Guo et al., 2018; Cohen et al., 2019; Levine & Feizi, 2020; Yang et al., 2020; Sulam et al., 2020).
@@
Besides data distribution, additional issues arise from the training algorithms. Vardi et al. (2022); Frei et al. (2023) have shown that when the data consists of clusters with small inter-cluster correlation, a shallow (two-layer) ReLU network trained by gradient flow generalizes well while failing to be robust against adversarial attacks of small radius, despite the existence of a much more robust classifier that can be explicitly constructed from a shallow network. This study unveils new challenges in the search for robust classifiers: Even if we know a robust classifier exists for certain data distribution, the implicit bias from our training algorithm (the choice of network architecture, optimization algorithm, etc.) may prevent us from finding it.
Paper contributions. In this paper, we show that under the same setting studied in Frei et al. (2023), the implicit bias of gradient flow that leads to non-robust networks can be altered to favor robust networks by modifying the ReLU activation. Specifically, we consider a data distribution consisting of a mixture of $K$ Gaussians, referred to as subclasses, which have small inter-subclass correlation and are grouped into two superclasses/classes. When training a shallow network to predict the class membership of a given data, we show, partially with formal theorems, and partially with conjectures validated by empirical experiments, that:
  • If the activation is a ReLU, neurons (rows of first-layer weight matrix) within the network tend to learn only the average direction of each class, leading to a classifier that generalizes well on clean data, but is vulnerable to an adversarial attack with $l_{2}$ radius $\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$, i.e. the trained network is non-robust with many subclasses. This leads to a new neural alignment perspective on the nonrobustness of ReLU networks identified by Frei et al. (2023).
@@
Our analysis (1) highlights the importance of the interplay between data structure and network architecture in determining the robustness of the trained network, (2) explains how the implicit bias (regularization) of training a ReLU network fails to exploit the data structure and leads to non-robust networks, and (3) shows how the issue is resolved by using a polynomial ReLU activation function. Moreover, numerical experiments on real datasets show that shallow networks with our generalized ReLU activation functions are much more robust than those with ReLU activation.
@@
Relation to existing analysis on implicit bias of neural networks. Our discussion is theory-centric. It builds upon the analysis of the implicit bias of training algorithms, but it is significantly different from prior theoretical analyses. The implicit bias of training algorithms has been studied extensively over past years for various architectures, including linear networks Saxe et al. (2014); Gunasekar et al. (2017); Ji & Telgarsky (2019); Woodworth et al. (2020); Min et al. (2021); Stöger & Soltanolkotabi (2021); Jacot et al. (2021); Wang & Jacot (2023), and nonlinear networks Lyu & Li (2019); Ji & Telgarsky (2020); Maennel et al. (2018); Chizat & Bach (2020); Boursier et al. (2022); Min et al. (2024); Wang & Ma (2023); Frei et al. (2022, 2023); Kumar & Haupt (2024); Abbe et al. (2023); Tarzanagh et al. (2023), and from different perspectives, including max-margin Lyu & Li (2019); Chizat & Bach (2020); Tarzanagh et al. (2023), min-norm Gunasekar et al. (2017); Min et al. (2021), sparsity/low-rankness Saxe et al. (2014); Woodworth et al. (2020); Wang & Jacot (2023); Abbe et al. (2023), and alignment Maennel et al. (2018); Kumar & Haupt (2024). While most works consider these implicit biases toward simple networks beneficial and prominent explanations for the success of neural networks in practice, and some existing work  (Faghri et al., 2021) has even shown that such biases help robustness in the cases of linear regression; few works Frei et al. (2023); Boursier & Flammarion (2024) discuss the potential harm caused by such biases. Our work takes one step further by proposing fixes to the harms identified by prior work (Frei et al., 2023), which sheds light on the potential of using deep learning theory to not only understand but to also improve neural networks in practice.
@@
Paper organization. Our main formal results regarding adversarial robustness will be shown (in Section 3) for explicitly constructed classifiers that can be realized by shallow networks with different activation functions. Then we connect these robustness results to those of actual trained shallow networks by a conjecture that shallow networks, when initialized properly, can learn such constructions via gradient flow training. The rationale behind this conjecture is carefully explained in Section 4, together with a preliminary theoretical analysis that supports this conjecture. Lastly, we empirically verify our conjecture in Section 5, followed by numerical experiments showing that our proposed new activation improves the robustness of shallow networks on real datasets.
@@
Notation: We denote the the inner product between vectors $\boldsymbol{x}$ and $\boldsymbol{y}$ by $\left\langle\boldsymbol{x},\boldsymbol{y}\right\rangle=\boldsymbol{x}^{\top}\boldsymbol{y}$, and the cosine of the angle between them as $\cos(\boldsymbol{x},\boldsymbol{y})=\langle\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|},\frac{\boldsymbol{y}}{\|\boldsymbol{y}\|}\rangle$. For an $n\times m$ matrix $\mathbf{A}$, we let $\|\boldsymbol{A}\|$ and $\|\boldsymbol{A}\|_{F}$ denote the spectral norm and Frobenius norm of $\boldsymbol{A}$, respectively. We also define $\mathbb{1}_{A}$ as the indicator for a statement $A$: $\mathbb{1}_{A}=1$ if $A$ is true and $\mathbb{1}_{A}=0$ otherwise. We also let $\mathcal{N}(\boldsymbol{\boldsymbol{\mu}},\boldsymbol{\Sigma}^{2})$ denote the normal distribution with mean $\boldsymbol{\boldsymbol{\mu}}$ and covariance matrix $\boldsymbol{\Sigma}^{2}$, and $\text{Unif}(S)$ denote the uniform distribution over a set $S$. Lastly, we let $[N]$ denote the integer set $\{1,\cdots,N\}$.
2 Problem Settings
@@
  • $(X\mid Z\!=\!k)\sim\mathcal{N}\left(\boldsymbol{\boldsymbol{\mu}}_{k},\frac{\alpha^{2}}{D}\boldsymbol{I}\right)$, $\forall 1\leq k\leq K$, where $\alpha>0$ is the intra-subclass variance, and $\{\boldsymbol{\mu}_{1},\cdots,\boldsymbol{\mu}_{K}\}$ are subclass centers that forms an orthonormal basis of a $K$-dimensional subspace in $\mathbb{R}^{D}$.
@@
Classification task and robustness of a classifier. The learning task we consider is to find a binary classifier $f(\boldsymbol{x}):\mathbb{R}^{D}\rightarrow\mathbb{R}$ such that given a randomly drawn $(\boldsymbol{x},y)$, $\mathrm{sign}(f(\boldsymbol{x}))$ predicts the correct class label $y$, i.e., $f(\boldsymbol{x})y>0$, with high probability. Moreover, we are interested in the robustness of $f(\boldsymbol{x})$ against additive adversarial attacks/perturbations on $\boldsymbol{x}$ with $l_{2}$-norm bounded by some constant $r>0$ (called the attack radius). Specifically, given a randomly drawn $(\boldsymbol{x},y)$ from $\mathcal{D}_{X,Y}$, we would like $\inf_{\|\boldsymbol{d}\|=1}f(\boldsymbol{x}+r\boldsymbol{d})y>0$ to hold with high probability (formal statement in the next section.) for $r$ as large as possible.
Gradient flow training. A candidate classifier can be realized from a neural network $f(\boldsymbol{x};\boldsymbol{\theta})$ parametrized by its networks weights $\boldsymbol{\theta}$. Given a size-$n$ training dataset $\{(\boldsymbol{x}_{i},y_{i}),i=1,\cdots,n\}$, where each training sample is randomly drawn from $\mathcal{D}_{X,Y}$, we define a loss function
@@
where $\ell(y,\hat{y}):\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ is either the exponential loss $\ell(y,\hat{y})=\exp\left(-y\hat{y}\right)$ or the logistic loss $\ell(y,\hat{y})=\log\left(1+\exp\left(-y\hat{y}\right)\right)$. We train the network using gradient flow (GF), $\dot{\boldsymbol{\theta}}(t)\in\partial\mathcal{L}(\boldsymbol{\theta}(t))$, where $\partial\mathcal{L}$ denotes the Clarke subdifferential (Clarke, 1990) w.r.t. $\boldsymbol{\theta}$. With proper initialization $\boldsymbol{\theta}(0)$, one expects the trained network weights $\boldsymbol{\theta}(T)$ (for some large $T>0$) to be close to some minimizer of $\mathcal{L}(\boldsymbol{\theta})$ and $f(\boldsymbol{x};\boldsymbol{\theta}(T))$ be a good classifier for $\mathcal{D}_{X,Y}$.
@@
that is, for each neuron $\boldsymbol{w}_{j}$, when compared to ReLU activation ($p=1$), the post-activation value is much smaller (penalized) if the angle separation between $\boldsymbol{x}$ and $\boldsymbol{w}_{j}$ is large. Such penalties, when $p>1$, promote the alignment between training data samples and neurons, effectively altering the implicit bias of the gradient flow training, and resulting in trained networks that are more robust. We further elaborate on this point in Section 4. In addition, such generalized ReLU activation keeps the function $f_{p}$ to be 2-positive-homogeneous w.r.t. its parameters $\boldsymbol{\theta}$, i.e., $f_{p}(\boldsymbol{x};\gamma\boldsymbol{\theta})=\gamma^{2}f_{p}(\boldsymbol{x};\boldsymbol{\theta})$ for any $\boldsymbol{\theta}$ and $\gamma>0$. Many existing analyses (Du et al., 2018; Lyu et al., 2021; Kumar & Haupt, 2024) on 2-positive-homogeneous networks apply to the generalized pReLU networks. Lastly, it is worth noting that the concept of promoting alignment between data points and neurons via cosine penalty has been explored for improving interpretability (Böhle et al., 2022) of the trained network.
@@ 3 Main Results on Adversarial Robustness
$F^{(p)}(\boldsymbol{x})=\sum\nolimits_{k=1}^{K_{1}}\sigma^{p}(\left\langle\boldsymbol{\mu}_{k},\boldsymbol{x}\right\rangle)-\sum\nolimits_{k=K_{1}+1}^{K}\sigma^{p}(\left\langle\boldsymbol{\mu}_{k},\boldsymbol{x}\right\rangle)\,$ (2)
@@
where $\bar{\boldsymbol{\mu}}_{+}=\frac{1}{\sqrt{K_{1}}}\sum_{k=1}^{K_{1}}\boldsymbol{\mu}_{k}$ and $\bar{\boldsymbol{\mu}}_{-}=\frac{1}{\sqrt{K_{2}}}\sum_{k=K_{1}+1}^{K}\boldsymbol{\mu}_{k}$ are, respectively, the average direction of the positive and negative subclass centers, with $\bar{\boldsymbol{\mu}}_{+}\!\in\!\mathbb{S}^{D-1}$ and $\bar{\boldsymbol{\mu}}_{-}\!\in\!\mathbb{S}^{D-1}$.
@@
Realizability of $F(\boldsymbol{x})$ and $F^{(p)}(\boldsymbol{x})$ by pReLU networks. We first notice that both $F(\boldsymbol{x})$ and $F^{(p)}(\boldsymbol{x})$ can be realized (non-uniquely) by $f_{p}(\boldsymbol{x};\{\boldsymbol{w}_{j},v_{j}\}_{j=1}^{h})$ with some choices for the weights $\{\boldsymbol{w}_{j},v_{j}\}_{j=1}^{h}$, stated as the following claim. (Verifying this claim is straightforward and we leave it to Appendix C)
@@
In other words, the classifier $F(\boldsymbol{x})$ corresponds to a vanilla shallow ReLU network whose non-zero neurons are either aligned with $\bar{\boldsymbol{\mu}}_{+}$, the average direction of positive subclass centers, or $\bar{\boldsymbol{\mu}}_{-}$, the negative counterpart. On the contrary, $F^{(p)}(\boldsymbol{x})$ corresponds to a pReLU network whose non-zero neurons are aligned with one of the subclass centers. As we will see soon, both $F(\boldsymbol{x})$ and $F^{(p)}(\boldsymbol{x})$ achieve high prediction accuracy on clean samples from $D_{X,Y}$, i.e., learning subclass centers is not necessary for good generalization, but $F^{(p)}(\boldsymbol{x})$ is much more robust than $F(\boldsymbol{x})$ against adversarial perturbation: learning subclass centers significantly improve robustness.
Generalization and Robustness of $F(\boldsymbol{x})$ and $F^{(p)}(\boldsymbol{x})$. Having established that both $F$ and $F^{(p)}$ are realizable by a pReLU network $f_{p}$, we now study the generalization performance and robustness of $F$ and $F^{(p)}$. Under the conjecture that both $F$ and $F^{(p)}$ can be learned by gradient flow training on a pReLU network $f_{p}$, we expect that such generalization and robustness properties to extend to $f_{p}$.
@@
Given classifiers $F(\boldsymbol{x})$, $F^{(p)}(\boldsymbol{x})$ defined in (3),(2), and a test sample $(\boldsymbol{x},y)\sim\mathcal{D}_{X,Y}$, we have, for some constant $C>0$,
@@
Given classifiers $F(\boldsymbol{x})$, $F^{(p)}(\boldsymbol{x})$ defined in (3),(2), and a test sample $(\boldsymbol{x},y)\sim\mathcal{D}_{X,Y}$, the following statement is true for the same constant $C$ in Proposition 1:
@@
We refer the readers to Appendix C for the proofs for Proposition 1 and Theorem 1. Our theoretical results should be understood in the high-dimensional or low-intra-class-variance regime so that $\frac{D}{\alpha^{2}}\gg K^{2}\log K$. On the one hand, Theorem 1 shows that $F(\boldsymbol{x})$, a classifier that corresponds to shallow ReLU networks whose neurons are aligned with either $\bar{\boldsymbol{\mu}}_{+}$ or $\bar{\boldsymbol{\mu}}_{-}$, is vulnerable to an adversarial attack of radius $\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$, which diminishes as the number of subclasses grows. On the other hand, Theorem 1 shows that $F^{(p)}(\boldsymbol{x})$, $p\geq 2$, a classifier that corresponds to shallow pReLU networks whose neurons are aligned with one of the subclass centers, is $\mathcal{O}\left(1\right)$-robust against adversarial attacks.
@@
Complementary results (in Appendix C.4) to Theorem 1 show that $F(\boldsymbol{x})$ is robust against any attack with $l_{2}$-norm slightly smaller than $\frac{1}{\sqrt{K}}$, and $F^{(p)}(\boldsymbol{x})$ is not robust to some attack with $l_{2}$-norm slightly larger than $\frac{\sqrt{2}}{2}$. Therefore, $\frac{1}{\sqrt{K}}$ (or $\frac{\sqrt{2}}{2}$, resp.) is the “critical" attack radius for $F(\boldsymbol{x})$ (or $F^{(p)}(\boldsymbol{x})$, resp.). We conjecture this is also the case for the associated trained networks, which we verify in Section 5.1.
@@
Given Conjecture 1, Theorem 1 can be used to infer the robustness of the networks $f_{p}(\boldsymbol{x};\boldsymbol{\theta}(T))$ obtained from gradient flow training with small initialization. When training a vanilla ReLU network in this regime, we expect the trained network to be close to $F(\boldsymbol{x})$, thus non-robust against $\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$-attacks. When training a pReLU network with $p\geq 3$, we expect the trained network to be close to $F^{(p)}(\boldsymbol{x})$, which is more robust than its ReLU counterpart.

### v2

@@ Abstract
The implicit bias of gradient-based training algorithms has been considered mostly beneficial as it leads to trained networks that often generalize well. However, Frei et al. (2023) show that such implicit bias can harm adversarial robustness. Specifically, they show that if the data consists of clusters with small inter-cluster correlation, a shallow (two-layer) ReLU network trained by gradient flow generalizes well, but it is not robust to adversarial attacks of small radius. Moreover, this phenomenon occurs despite the existence of a much more robust classifier that can be explicitly constructed from a shallow network. In this paper, we extend recent analyses of neuron alignment to show that a shallow network with a polynomial ReLU activation (pReLU) trained by gradient flow not only generalizes well but is also robust to adversarial attacks. Our results highlight the importance of the interplay between data structure and architecture design in the implicit bias and robustness of trained networks.
@@ 1 Introduction
Behind the success of deep neural networks in many application domains lies their vulnerability to adversarial attacks, i.e., small and human-imperceptible perturbations to the input data. Such a phenomenon was observed in the seminal paper of Szegedy et al. (2014) and has motivated a large body of work on building defenses against such attacks (Shafahi et al., 2019; Papernot et al., 2016; Wong et al., 2019; Guo et al., 2018; Cohen et al., 2019; Levine & Feizi, 2020; Yang et al., 2020; Sulam et al., 2020; Kinfu & Vidal, 2022).
@@
Besides data distribution, additional issues arise from the training algorithms. For example, Vardi et al. (2022) and Frei et al. (2023) show that when the data consists of clusters with small inter-cluster correlation, a shallow (two-layer) ReLU network trained by gradient flow generalizes well but fails to be robust against adversarial attacks of small radius, despite the existence of a much more robust classifier that can be explicitly constructed from a shallow network. This study unveils new challenges in the search for robust classifiers: Even if we know robust classifiers exist for certain data distribution, the implicit bias from our training algorithm (the choice of network architecture, optimization algorithm, etc.) may prevent us from finding it.
Figure 1: Visualizing the training of a pReLU network under small initialization (Details explained in later sections). The dataset has its positive class sampled from two subclasses. (a) At initialization, all neurons have small norms and point toward random directions; When $p=1$ (vanilla ReLU network), (b) During the alignment phase, the neuron directions are aligned with either of the average class centers $\bar{\boldsymbol{\mu}}_{+}$ and $\bar{\boldsymbol{\mu}}_{-}$; (c) During the second phase, neurons keep the alignment with $\bar{\boldsymbol{\mu}}_{+}$ and $\bar{\boldsymbol{\mu}}_{-}$ while growing their norms; When $p=3$, (d) neurons learn subclass centers during alignment phase and (e) keep the alignment in the second phase. Note: the neurons pointing toward directions other than class/subclass centers are not activated by any data point and have small norms throughout training.
Paper contributions. In this paper, we show that under the same setting studied in Frei et al. (2023), the implicit bias of gradient flow that leads to non-robust networks can be altered to favor robust networks by modifying the ReLU activation. Specifically, we consider a data distribution consisting of a mixture of $K$ Gaussians, referred to as subclasses, which have small inter-subclass correlation and are grouped into two superclasses/classes. When training a two-layer binary classification network, we show, with formal theorems, and with conjectures validated experimentally, that (also illustrated in Figure 1):
  • If the activation is a ReLU, neurons (rows of the first layer weight matrix) tend to learn only the average direction of each class, leading to a classifier that generalizes well on clean data, but is vulnerable to an adversarial attack with $l_{2}$ radius $\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$, i.e. the trained network is non-robust with many subclasses. This leads to a new neural alignment perspective on the nonrobustness of ReLU networks identified by Frei et al. (2023).
@@
Our analysis (1) highlights the importance of the interplay between data structure and network architecture in determining the robustness of the trained network, (2) explains how the implicit bias (regularization) of training a ReLU network fails to exploit the data structure and leads to non-robust networks, and (3) shows how the issue is resolved by using a polynomial ReLU activation function. Moreover, numerical experiments on real datasets show that shallow networks with our generalized ReLU activation functions are much more robust than those with a ReLU activation.
@@
Relation to existing analysis on implicit bias of neural networks. Our discussion is theory-centric. It builds upon the analysis of the implicit bias of training algorithms, but it is significantly different from prior theoretical analyses. The implicit bias of training algorithms has been studied extensively over past years for various architectures, including both linear networks Saxe et al. (2014); Gunasekar et al. (2017); Ji & Telgarsky (2019); Woodworth et al. (2020); Min et al. (2021); Stöger & Soltanolkotabi (2021); Jacot et al. (2021); Wang & Jacot (2023), and nonlinear networks Lyu & Li (2019); Ji & Telgarsky (2020); Maennel et al. (2018); Chizat & Bach (2020); Boursier et al. (2022); Min et al. (2024); Wang & Ma (2023); Frei et al. (2022, 2023); Kumar & Haupt (2024); Abbe et al. (2023); Tarzanagh et al. (2023). Moreover, this phenomenon has been studied from different perspectives, including max-margin Lyu & Li (2019); Chizat & Bach (2020); Tarzanagh et al. (2023), min-norm Gunasekar et al. (2017); Min et al. (2021), sparsity/low-rankness Saxe et al. (2014); Woodworth et al. (2020); Wang & Jacot (2023); Abbe et al. (2023), and alignment Maennel et al. (2018); Min et al. (2024); Kumar & Haupt (2024). While most works consider these implicit biases beneficial for the success of neural networks in practice, and some existing work  (Faghri et al., 2021) has even shown that such biases help robustness in the cases of linear regression; few works Frei et al. (2023); Boursier & Flammarion (2024) discuss the potential harm caused by such biases. Our work takes one step further by proposing fixes to the harms identified by prior work (Frei et al., 2023), which sheds light on the potential of using deep learning theory to not only understand but also improve neural networks in practice.
@@
Paper organization. Our main formal results regarding adversarial robustness will be shown (in Section 3) for explicitly constructed classifiers that can be realized by shallow networks with different activation functions. Then we connect these robustness results to those of actual trained shallow networks by a conjecture that shallow networks, when initialized properly, can learn such constructions via gradient flow training. The rationale behind this conjecture is carefully explained in Section 4, together with a preliminary theoretical analysis that supports this conjecture. Lastly, we empirically verify our conjecture in Section 5 via experiments on synthetic data, and then we conduct numerical experiments on real datasets, showing that our proposed new activation improves the robustness of shallow networks.
@@
Notation: We denote the the inner product between vectors $\boldsymbol{x}$ and $\boldsymbol{y}$ by $\left\langle\boldsymbol{x},\boldsymbol{y}\right\rangle=\boldsymbol{x}^{\top}\boldsymbol{y}$, and the cosine of the angle between them as $\cos(\boldsymbol{x},\boldsymbol{y})=\langle\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|},\frac{\boldsymbol{y}}{\|\boldsymbol{y}\|}\rangle$. For an $n\times m$ matrix $\mathbf{A}$, we let $\|\boldsymbol{A}\|$ and $\|\boldsymbol{A}\|_{F}$ denote the spectral and Frobenius norm of $\boldsymbol{A}$, respectively. We also define $\mathbb{1}_{A}$ as the indicator for a statement $A$: $\mathbb{1}_{A}=1$ if $A$ is true and $\mathbb{1}_{A}=0$ otherwise. We also let $\mathcal{N}(\boldsymbol{\boldsymbol{\mu}},\boldsymbol{\Sigma}^{2})$ denote the normal distribution with mean $\boldsymbol{\boldsymbol{\mu}}$ and covariance matrix $\boldsymbol{\Sigma}^{2}$, and $\text{Unif}(S)$ denote the uniform distribution over a set $S$. Lastly, we let $[N]$ denote the integer set $\{1,\cdots,N\}$.
2 Problem Setting
@@
  • $(X\mid Z\!=\!k)\sim\mathcal{N}\left(\boldsymbol{\boldsymbol{\mu}}_{k},\frac{\alpha^{2}}{D}\boldsymbol{I}\right)$, $\forall 1\leq k\leq K$, where $\alpha>0$ is the intra-subclass variance, and $\{\boldsymbol{\mu}_{1},\cdots,\boldsymbol{\mu}_{K}\}$ are subclass centers that forms an orthonormal basis of a $K$-dimensional subspace in $\mathbb{R}^{D}$, where $K<D$.
@@
Classification task and robustness of a classifier. The learning task we consider is to find a binary classifier $f(\boldsymbol{x}):\mathbb{R}^{D}\rightarrow\mathbb{R}$ such that given a randomly drawn $(\boldsymbol{x},y)$, $\mathrm{sign}(f(\boldsymbol{x}))$ predicts the correct class label $y$, i.e., $f(\boldsymbol{x})y>0$, with high probability. Moreover, we are interested in the robustness of $f(\boldsymbol{x})$ against additive adversarial attacks/perturbations on $\boldsymbol{x}$ with $l_{2}$-norm bounded by some constant $r>0$ (called the attack radius). Specifically, given a randomly drawn $(\boldsymbol{x},y)$ from $\mathcal{D}_{X,Y}$, we would like $\inf_{\|\boldsymbol{d}\|=1}f(\boldsymbol{x}+r\boldsymbol{d})y>0$ to hold with high probability (formal statement in Section 3) for $r$ as large as possible.
Gradient flow training. A candidate classifier can be realized from a neural network $f(\boldsymbol{x};\boldsymbol{\theta})$ parameterized by its network weights $\boldsymbol{\theta}$. Given a size-$n$ training dataset $\{(\boldsymbol{x}_{i},y_{i}),i=1,\cdots,n\}$, where each training sample is randomly drawn from $\mathcal{D}_{X,Y}$, we define a loss function
@@
where $\ell(y,\hat{y}):\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ is either the exponential loss $\ell(y,\hat{y})=\exp\left(-y\hat{y}\right)$ or the logistic loss $\ell(y,\hat{y})=\log\left(1+\exp\left(-y\hat{y}\right)\right)$. We train the network using gradient flow (GF), $\dot{\boldsymbol{\theta}}(t)\in\partial\mathcal{L}(\boldsymbol{\theta}(t))$, where $\partial\mathcal{L}$ denotes the Clarke subdifferential (Clarke, 1990) w.r.t. $\boldsymbol{\theta}$. With proper initialization $\boldsymbol{\theta}(0)$, one expects the trained network weights $\boldsymbol{\theta}(T)$ (for some large $T>0$) to be close to some minimizer of $\mathcal{L}(\boldsymbol{\theta})$ and $f(\boldsymbol{x};\boldsymbol{\theta}(T))$ to be a good classifier for $\mathcal{D}_{X,Y}$.
@@
that is, for each neuron $\boldsymbol{w}_{j}$, when compared to ReLU activation ($p=1$), the post-activation value is much smaller (penalized) if the angle separation between $\boldsymbol{x}$ and $\boldsymbol{w}_{j}$ is large. When $p>1$, such penalties, as we will see later, promote the alignment between training data samples and neurons, and result in trained networks that capture well the intrinsic structure of the data. In addition, such generalized ReLU activation makes the function $f_{p}$ positively homogeneous of degree two w.r.t. its parameters $\boldsymbol{\theta}$, i.e., $f_{p}(\boldsymbol{x};\gamma\boldsymbol{\theta})=\gamma^{2}f_{p}(\boldsymbol{x};\boldsymbol{\theta})$ for any $\boldsymbol{\theta}$ and $\gamma>0$. Many existing analyses (Du et al., 2018; Lyu et al., 2021; Kumar & Haupt, 2024) on positively homogeneous networks apply to the generalized pReLU networks. Lastly, it is worth noting that the concept of promoting alignment between data points and neurons via cosine penalty has been explored for improving interpretability (Böhle et al., 2022) of the trained network.
@@ 3 Main Results on Adversarial Robustness
$F^{(p)}(\boldsymbol{x})=\sum_{k=1}^{K_{1}}\sigma^{p}(\left\langle\boldsymbol{\mu}_{k},\boldsymbol{x}\right\rangle)-\sum_{k=K_{1}+1}^{K}\sigma^{p}(\left\langle\boldsymbol{\mu}_{k},\boldsymbol{x}\right\rangle)\,$ (2)
@@
where $\bar{\boldsymbol{\mu}}_{+}\!=\!\frac{1}{\sqrt{K_{1}}}\sum_{k=1}^{K_{1}}\boldsymbol{\mu}_{k}$ and $\bar{\boldsymbol{\mu}}_{-}\!=\!\frac{1}{\sqrt{K_{2}}}\sum_{k=K_{1}+1}^{K}\boldsymbol{\mu}_{k}$ are, respectively, the average direction of the positive and negative subclass centers. Here the average direction is computed by taking the sum of all $\boldsymbol{\mu}_{k}$s, then normalize it, thus we have $\bar{\boldsymbol{\mu}}_{+}\!\in\!\mathbb{S}^{D-1}$ and $\bar{\boldsymbol{\mu}}_{-}\!\in\!\mathbb{S}^{D-1}$.
@@
Realizability of $F(\boldsymbol{x})$ and $F^{(p)}(\boldsymbol{x})$ by pReLU networks. We first notice that both $F(\boldsymbol{x})$ and $F^{(p)}(\boldsymbol{x})$ can be realized (non-uniquely) by a pReLU network $f_{p}(\boldsymbol{x};\{\boldsymbol{w}_{j},v_{j}\}_{j=1}^{h})$ for suitable choices of the weights $\{\boldsymbol{w}_{j},v_{j}\}_{j=1}^{h}$, as stated in the following claim. (Verifying this claim is straightforward and we leave it to Appendix C).
@@
Notice that the classifier $F$ corresponds to a vanilla shallow ReLU network whose non-zero neurons are either aligned with $\bar{\boldsymbol{\mu}}_{+}$, the average direction of the positive subclass centers, or $\bar{\boldsymbol{\mu}}_{-}$, the negative counterpart. On the other hand, $F^{(p)}$ corresponds to a pReLU network whose non-zero neurons are aligned with one of the subclass centers. As we will see soon, both $F(\boldsymbol{x})$ and $F^{(p)}(\boldsymbol{x})$ achieve high prediction accuracy on clean samples $(\boldsymbol{x},y)$ from $D_{X,Y}$, i.e., learning subclass centers is not necessary for good generalization. However, we will also show that $F^{(p)}$ is much more robust against adversarial perturbation than $F$. This is why we argue that learning subclass centers significantly improve robustness.
Generalization and Robustness of $F(\boldsymbol{x})$ and $F^{(p)}(\boldsymbol{x})$. Having established that both $F$ and $F^{(p)}$ are realizable by a pReLU network $f_{p}$, we now study the generalization performance and robustness of $F$ and $F^{(p)}$. Under the conjecture that both $F$ and $F^{(p)}$ can be learned by gradient flow training on a pReLU network $f_{p}$, we expect that such generalization and robustness properties extend to $f_{p}$.
@@
Given a test sample $(\boldsymbol{x},y)\sim\mathcal{D}_{X,Y}$, and classifiers $F$ and $F^{(p)}$, $p\geq 1$, defined in (3) and (2), resp., there is a constant $C$ such that
@@
$\mathbb{P}\left(F^{(p)}(\boldsymbol{x})y>0\right)$ $\geq 1-2(K+1)\exp\left(-\frac{CD}{\alpha^{2}K^{2}}\right).$
@@
Given a test sample $(\boldsymbol{x},y)\sim\mathcal{D}_{X,Y}$, and classifiers $F$ and $F^{(p)}$, defined in (3) and (2), resp., the following statement is true for the same constant $C$ in Proposition 1:
@@
We refer the reader to Appendix C for the proofs for Proposition 1 and Theorem 1. Our theoretical results should be understood in the high-dimensional or low-intra-class-variance regime so that $\frac{D}{\alpha^{2}}\gg K^{2}\log K$. On the one hand, Theorem 1 shows that $F$, a classifier that corresponds to shallow ReLU networks whose neurons are aligned with either $\bar{\boldsymbol{\mu}}_{+}$ or $\bar{\boldsymbol{\mu}}_{-}$, is vulnerable to an adversarial attack of radius $\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$, which diminishes as the number of subclasses grows. On the other hand, Theorem 1 shows that $F^{(p)}$, $p\geq 2$, a classifier that corresponds to shallow pReLU networks whose neurons are aligned with one of the subclass centers, is $\mathcal{O}\left(1\right)$-robust against adversarial attacks.
@@
Complementary results to Theorem 1 in Appendix C.4 show that $F$ is robust against any attack with $l_{2}$-norm slightly smaller than $\frac{1}{\sqrt{K}}$, and $F^{(p)}$ is not robust to some attack with $l_{2}$-norm slightly larger than $\frac{\sqrt{2}}{2}$. Therefore, $\frac{1}{\sqrt{K}}$ and $\frac{\sqrt{2}}{2}$ are the “critical" attack radius for $F$ and $F^{(p)}$, resp. We conjecture this is also the case for the associated trained networks, which we verify in Section 5.1.
@@
Given Conjecture 1, Theorem 1 can be used to infer the robustness of the networks $f_{p}(\boldsymbol{x};\boldsymbol{\theta}(T))$ obtained from gradient flow training with small initialization. When training a vanilla ReLU network in this regime, we expect the trained network to be close to $F$, thus non-robust against $\mathcal{O}\left(\frac{1}{\sqrt{K}}\right)$-attacks. When training a pReLU network with $p\geq 3$, we expect the trained network to be close to $F^{(p)}$, which is more robust than its ReLU counterpart.